OpenStudy (annie96):

How many solutions are there to the following system of equations? 4x-14y=6 -2x+7y=-3 0 1 2 infinitely many

OpenStudy (anonymous):

Note that \(\large 4x-14y = 6 \implies -2(-2x+7y) = -2(-3) \implies -2x+7y = -3\). With that said, can you figure out how many solutions you have now? :-)

OpenStudy (anonymous):

Two equations are constant multiple of each other. In order to solve for two variables x and y you need two unique equations.

OpenStudy (anonymous):

It have 0 Solutions since both the equations are same :)

OpenStudy (annie96):

Thank you.

OpenStudy (annie96):

Help.

OpenStudy (anonymous):

Since the system of equations reduces to one line \(\large -2x+7y = -3\), then there are an infinite number of solutions because there are an infinite number of points \(\large x,y\) that lie on tje line. Does this clarify things? :-)

OpenStudy (annie96):

yes thank you.